OpenStudy (anonymous):

18x+6y=19 -3x=y+6 solve system of equations by substitution method

OpenStudy (anonymous):

18x+6y=19 -3x=y+6 ====> y = -3x - 6 18x+6y=19 18x+6(-3x - 6) =19 finish it